OpenStudy (anonymous):

Find all Vertical Asymptotes. The answer is x=1;x=-2, but I can't figure out how. The x^3 in the denominator threw everything off.

OpenStudy (anonymous):

\[h(x)=\frac{ x^3+x^2+x+1 }{ x^3-3x+2 }\]

OpenStudy (anonymous):

To find vertical asymptotes you are generally concerned with the denominator. I like to factor the denominator to make the job easier:\[x ^{2}-3x+2=(x-1)(x-2)\] Now set that equal to zero and you obtain that the function is undefined (a vertical asymptote) when x=1 and x-2.

OpenStudy (anonymous):

That should say x=2 not x-2. Sorry typo ;p

OpenStudy (anonymous):

However, in the denominator its \[x^3\]not\[x^2\]

OpenStudy (anonymous):

Oh I'm sorry I thought that was a 2. You're going to make me work a little harder to factor that lol. Just a sec.

OpenStudy (anonymous):

So the correct denominator is \[x ^{3}-3x+2=(x+2)(x-1)^{2}=0\] From this we can tell that the roots are x=-2 and x=1.

OpenStudy (anonymous):

Oh so that's how you factor it.

OpenStudy (anonymous):

Do you know how to use synthetic division? I prefer that method when dividing large polynomials by potential factors.

OpenStudy (anonymous):

I said large. What I meant was "any." lol ;p

OpenStudy (anonymous):

I see. Well Thank you for helping me Jonathan.

OpenStudy (anonymous):

Np good luck with your classes.

OpenStudy (anonymous):

Thank you.